OpenStudy (anonymous):

Use the simple interest formula to find the missing value. Rates are annual rates. Assume 360 days per year and 30 days per month. p=$6404, r=?,t=4 months, i=$202.79

OpenStudy (helder_edwin):

if r is the annual interest rate then r/12 is the monthly interest rate

OpenStudy (helder_edwin):

so u have \[ \large i=p\cdot\frac{4r}{12} \]

OpenStudy (helder_edwin):

do u get this?

OpenStudy (anonymous):

kinda.. my answer choices are A.r=10.53% B.r=0.10% C.r=7.4% D.r=9.5% so would the answer be A ..?

OpenStudy (helder_edwin):

u have i=202.79 p=6404 plug in and solve for r it is easy

OpenStudy (anonymous):

how do i plug it in?

OpenStudy (anonymous):

9.5?!

OpenStudy (helder_edwin):

OK \[ \large 202.79=6404\cdot\frac{4r}{12} \]

OpenStudy (helder_edwin):

this is an equation for r can u solve this?

OpenStudy (helder_edwin):

u r right it is 9.5%

OpenStudy (anonymous):

thankyou!!

OpenStudy (helder_edwin):

u r w